OpenStudy (anonymous):

Solve for x by completing the square of f(x)=2x^2+7x+3 Help please i have part but the rest is so confusing!

OpenStudy (anonymous):

sometimes it is just easier to use the quadratic formula, which is completing the square in one step

OpenStudy (mathstudent55):

You need to solve \(2x^2 + 7x + 3 = 0\) by completing the square?

OpenStudy (anonymous):

yes i do @mathstudent55

OpenStudy (anonymous):

i already did that part @satellite73, i have to do this part too.

OpenStudy (mathstudent55):

Ok. I'll tell you what the steps are. You do them and show me. I will go over them with you to make sure you're doing them correctly.

OpenStudy (mathstudent55):

The first step is that you must have an equation of the form \(x^2 + bx + c = 0\). The coefficient of the \(x^2\) term must be 1, so divide both sides of the equation by 2.

OpenStudy (anonymous):

no i was told by my teacher that completing the square was done differently @mathstudent55

OpenStudy (mathstudent55):

This is how I do it. \(2x^2+7x+3 = 0\) Divide the entire equation by 2: \(x^2+\dfrac{7}{2}x+\dfrac{3} {2} = 0\)

OpenStudy (anonymous):

she told me i had to subtract 3 from both sides

OpenStudy (mathstudent55):

Now subtract the constant term to the right side: \(x^2+\dfrac{7}{2}x ~~~~~~~= -\dfrac{3} {2} \)

OpenStudy (mathstudent55):

Ok. After you subtract the 3, then you divide both sides by 2?

OpenStudy (anonymous):

so it will be x^2+3.5=-3/2?

OpenStudy (anonymous):

@mathstudent55

OpenStudy (mathstudent55):

Exactly. That's what I have but with fractions.

OpenStudy (mathstudent55):

Only you meant 3.5x, not just 3.5 on the left side.

OpenStudy (anonymous):

right thought i typed the x. sorry, what do i do next? @mathstudent55

OpenStudy (mathstudent55):

When you write the last equation, leave a space between the 3.5x and the = sign. That's where we will add the complete the square term. \(x^2+3.5x~~~~~~~=-1.5\)

OpenStudy (anonymous):

..okay so whats the next step? @mathstudent55

OpenStudy (mathstudent55):

Let's go back to fractions. \(x^2+\dfrac{7}{2}x ~~~~~~~= -\dfrac{3} {2}\) The next step is the "complete the square step." We take half of the coefficient of x, in this case 7/2, and we square it. \(\left( \dfrac{7}{2} \right) = \dfrac{49}{4} \)

OpenStudy (mathstudent55):

Then we add that to both sides. \(x^2+\dfrac{7}{2}x +\color{red}{\dfrac{49}{4} }= -\dfrac{3} {2} +\color{red}{\dfrac{49}{4} }\)

OpenStudy (anonymous):

okay so now what? @mathstudent55

OpenStudy (mathstudent55):

Wait. Let's go back a few steps. Let me write it again.

OpenStudy (mathstudent55):

Let's go back to fractions. \(x^2+\dfrac{7}{2}x ~~~~~ =−\dfrac{3}{2}\). The next step is the "complete the square step." We take half of the coefficient of x, in this case \(\dfrac{7}{2}\).

OpenStudy (anonymous):

Okay now I'm super confused @mathstudent55

OpenStudy (mathstudent55):

\(\dfrac{7}{2}\) is the coefficient of x. We take half of it: \(\dfrac{7}{4} \) Now we square that: \(\dfrac{49}{16} \)

OpenStudy (anonymous):

how is half of 7/2 7/4? @mathstudent55

OpenStudy (mathstudent55):

Ok. I am sorry. I wrote above what to do but did it incorrectly. We need to take the coefficient of the x term. It's 7/2.

OpenStudy (mathstudent55):

The we need to take half of it. Half of 7/2 is 7/4.

OpenStudy (anonymous):

okay...? @mathstudent55 so instead of 49/4 its 49/16?

OpenStudy (mathstudent55):

This is why: \(\dfrac{7}{2} \div 2 = \dfrac{7}{2} \div \dfrac{2}{1} = \dfrac{7}{2} \times \dfrac{1}{2} = \dfrac{7}{4}\)

OpenStudy (mathstudent55):

Correct.

OpenStudy (mathstudent55):

Now we square 7/4 to get 49/16.

OpenStudy (mathstudent55):

That's what we add to both sides.

OpenStudy (mathstudent55):

We had this. \(x^2+\dfrac{7}{2}x ~~~~~ =−\dfrac{3}{2}\) Now we add 49/16 to both sides: \(x^2+\dfrac{7}{2}x + \dfrac{49}{16} =−\dfrac{3}{2} + \dfrac{49}{16}\)

OpenStudy (anonymous):

okay

OpenStudy (mathstudent55):

Since this whole procedure is completing the square on the left side, the left side now is the square of a binomial, so we write the left side as \((x + \dfrac{7}{4})^2 \).

OpenStudy (mathstudent55):

\(\left(x+\dfrac{7}{4} \right)^2 =−\dfrac{3}{2} + \dfrac{49}{16} \) Now we add the fractions on the right side. We need to use the LCD of 16.

OpenStudy (anonymous):

now i'm confused

OpenStudy (anonymous):

and i have a serious time limit i have to go teach bible school in 15 minutes

OpenStudy (anonymous):

@mathstudent55

OpenStudy (mathstudent55):

We're almost done. Ask your question. Where are you confused?

OpenStudy (anonymous):

how does the left side go from x^2+7/2x to a 7/4?

OpenStudy (anonymous):

where did the 49/16 go?

OpenStudy (anonymous):

@mathstudent55

OpenStudy (mathstudent55):

The entire idea behind the "complete the square" method is to turn the left side into a perfect square trinomial.

OpenStudy (anonymous):

so do the 7/2x +49/16 =7/4? @mathstudent55

OpenStudy (mathstudent55):

Remember the pattern: \((a + b)^2 = a^2 + 2ab + b^2\) We are doing it in reverse.

OpenStudy (anonymous):

but am i correct?@math&ing001

OpenStudy (mathstudent55):

We took \(a^2 + 2ab\). Then we figure out what we needed to add to get the \(b^2\) part. We added it. Then we went from \(a^2 + 2ab + b^2 \) to \((a + b)^2\).

OpenStudy (anonymous):

okay no what now? @mathstudent55

OpenStudy (mathstudent55):

We applied the pattern to do this to the left side. \(x^2 + \dfrac{7}{2}x + \dfrac{49}{16} = \left(x+\dfrac{7}{4} \right)^2 \)

OpenStudy (anonymous):

yes i know, how do i solve for x? from here? @mathstudent55

OpenStudy (mathstudent55):

Now we add the fractions on the right side suing a common denominator. \(\left(x+\dfrac{7}{4} \right)^2 =−\dfrac{3}{2} \times \dfrac{8}{8} + \dfrac{49}{16} \) \(\left(x+\dfrac{7}{4} \right)^2 =−\dfrac{24}{16} + \dfrac{49}{16} \) \(\left(x+\dfrac{7}{4} \right)^2 =\dfrac{25}{16} \) Now there is a principle that if x^2 = k (k is a number), then \(x = \sqrt{k}\) or \(x = -\sqrt{k} \). We apply the principle to our situation: \(x+\dfrac{7}{4} =\sqrt{\dfrac{25}{16}} \) or \(x+\dfrac{7}{4} =-\sqrt{\dfrac{25}{16}} \) \(x+\dfrac{7}{4} =\dfrac{5}{4} \) or \(x+\dfrac{7}{4} =-\dfrac{5}{4} \) \(x =-\dfrac{2}{4} \) or \(x =-\dfrac{12}{4} \) \(x =-\dfrac{1}{2} \) or \(x =-3 \)

OpenStudy (anonymous):

THANK YOU SO MUCH! @mathstudent55

OpenStudy (mathstudent55):

You're welcome.